# Dialectal Coverage And Generalization in Arabic Speech Recognition

### v2

@@ Abstract
Developing robust automatic speech recognition (ASR) systems for Arabic, a language characterized by its rich dialectal diversity and often considered a low-resource language in speech technology, demands effective strategies to manage its complexity. This study explores three critical factors influencing ASR performance: the role of dialectal coverage in pre-training, the effectiveness of dialect-specific fine-tuning compared to a multi-dialectal approach, and the ability to generalize to unseen dialects. Through extensive experiments across different dialect combinations, our findings offer key insights towards advancing the development of ASR systems [^1] The code with pre-trained & finetuned model weights is available on GitHub: https://github.com/mbzuai-nlp/ArTST for pluricentric languages like Arabic.
@@ 1 Introduction
The advent of large self-supervised acoustic models has transformed speech technology, enabling transfer learning and improving performance for both high-resource and low-resource languages. Prominent examples of such models include various versions of wav2vec Schneider et al. (2019); Baevski et al. (2020), HuBERT Hsu et al. (2021), and SpeechT5 Ao et al. (2021), which have predominantly been trained on English datasets. Their multi-lingual variants, e.g. XLS-R Babu et al. (2021) with 53 and 128 languages, in addition to many models that include both self-supervised and supervised pre-training, such as Whisper Radford et al. (2023) with approximately hundred supported languages, MMS Pratap et al. (2024) with thousands of languages, and UniSpeech Wang et al. (2021), underscore the potential for cross-lingual transfer learning for more inclusive ASR. Yet, while these models indeed show great potential for transfer learning to new languages, even those unseen in training Huang et al. (2013), they remain sub-optimal for specific target languages. A case in point is the Arabic Text and Speech Transformer (ArTST), a model pre-trained exclusively on Arabic, which has demonstrated superior performance for Modern Standard Arabic (MSA), surpassing larger multilingual models like Whisper and MMS in benchmark tests, in addition to establishing a new state-of-the-art (SOTA) performance compared to previous efforts for Arabic ASR. This highlights the advantage of monolingual pre-training when large amounts of unlabeled data for the target language are available. While the model showed some potential for dialectal coverage, it was trained and validated mainly on MSA data, which questions its applicability for spoken dialectal variants of Arabic. Arabic is a pluricentric language Schuppler et al. (2024), diverse in regional variations, and models trained on MSA frequently struggle to adapt to these variations. This limitation is particularly acute given that many Arabic dialects are underrepresented or considered low-resource in speech technology research. Consequently, there is a need for optimized ASR systems that embrace, rather than overlook, the linguistic diversity of the Arabic-speaking world.
@@
In light of these challenges, we conduct various investigations aimed at understanding and enhancing the dialectal diversity and performance of Arabic ASR systems. We focus on three inquiries aimed at optimizing potential strategies for integrating dialectal variation into ASR systems. First, we measure the impact of incorporating a broad collection of Arabic dialects during the model’s pre-training phase. We hypothesize that a wider dialectal foundation could improve the model’s performance across various dialects in the fine-tuning stage. Second, we quantify the comparative effectiveness of dialect-specific fine-tuning versus a more holistic, multi-dialectal fine-tuning strategy. The third question examines the model’s capacity for zero-shot transfer to dialects not explicitly included in fine-tuning. Our key findings from experiments spanning over 12 Arabic dialects are: (1) pre-training with more data and wider dialectal coverage improves performance across most dialectal variants, including MSA, (2) multi-dialectal fine-tuning improves performance for low-resource dialects, but may not be optimal for high-resource dialects, and (3) multi-dialectal pre-training and fine-tuning has higher potential for zero-shot transfer to unseen dialects.
Our pre-training checkpoints and joint models are trained exclusively on open-source data and will be released as open-source models.
@@ 2 Related Work
Recent research in Arabic speech recognition, as presented in Hussein et al. (2022), demonstrates the potential of contemporary deep learning techniques in decoding Arabic spoken language. However, this initial success was confined to Modern Standard Arabic (MSA), the formal variant predominant in news broadcasts and official communications. Large-scale multi-lingual ASR models, Whisper Radford et al. (2023) and MMS Pratap et al. (2024), cover many languages within their scope, including Arabic. They utilize language embeddings or adapters to enhance language coverage and performance within the same model, but their performance across languages vary considerably. Toyin et al. (2023) demonstrated state-of-the-art performance in multi-task training for Arabic speech recognition and synthesis, improving over much larger multi-lingual models, but their model was still trained predominantly on MSA. Their findings, however, demonstrate the model’s preliminary potential for processing Arabic dialects. Waheed et al. (2023) explored dialectal fine-tuning of large speech models like Whisper Radford et al. (2023), and XLS-R Babu et al. (2021), demonstrating their effectiveness on Egyptian and Moroccan dialects. The generalizability of these models to other Arabic dialects remains an open question.
3 Research Methodology
Our objective is to optimize state-of-the-art ASR models for dialectal diversity. To that end, we start with the current state-of-the-art model for Arabic, ArTST Toyin et al. (2023), as the foundation model for our investigation. Figure 1 illustrates the high-level architecture of ArTST for self-supervised pre-training and fine-tuning. This model is based on the SpeechT5 approach Ao et al. (2021), and supports multi-modal fine-tuning. The model was pre-trained on the MGB2 Ali et al. (2016) dataset, which consists of newswire data, mainly in MSA, with a small subset of dialectal variants. The model showed some potential for dialectal coverage in ASR, in addition to demonstrating SOTA performance in the task of dialect identification Ali et al. (2019). In this work, we focus on ASR performance across dialects, and attempt to quantify the effect of various design decisions in the development of such ASR models. To understand the factors that enable wider dialectal coverage, we explore the following questions:
@@ 3 Research Methodology
    Is pre-training on dialectal data essential for improving down-stream performance?
@@
    Is it better to fine-tune ASR models on multiple dialects or fine-tune only on the target dialect?
@@
    What is the effect of data augmentation on dialectal fine-tuning performance?
The following sections describe our experimental settings, results, and findings for each of these empirical questions.
Dataset Dialect Hours Words v1 v2
ASC MSA 3.7 hrs 20.58 K
CommVoice mix 133.24 hrs 494.83 K
ClArTTS CA 12 hrs 76.31 K
MGB2 mostly MSA 1000 hrs 7.31 M
QASR mostly MSA 2000 hrs 13.33 M
MGB3[ASR] EGY 2.83 hrs 18.93 K
MGB3[ADI] EGY 11.15 hrs
GLF 8.92 hrs
LAV 9.27 hrs
MSA 9.39 hrs
NOR 9.49 hrs
MGB5[ADI] ALG 115.7hrs
EGY 451.1 hrs
IRA 815.8 hrs
JOR 25.9 hrs
KSA 186.1 hrs
KUW 108.2 hrs
LEB 116.8 hrs
LIB 127.4 hrs
MAU 456.4 hrs
MOR 57.8 hrs
OMA 58.5 hrs
PAL 121.4 hrs
QAT 62.3 hrs
SUD 47.7 hrs
SYR 119.5 hrs
UAE 108.4 hrs
YEM 53.4 hrs
MADAR MOR ALG TUN 532.37K
LIB EGY LEV
IRA GLF YEM
NADI ALG BAH EGY 702.67K
IRA JOR KUW
LEB LIB MOR
OMN PAL QAT
SAU SUD SYR
TUN UAE YEM
Table 1: Summary of Dataset Statistics for Pre-Training: Hours of Audio, Word Counts, and Associated Dialects.
@@ 3.2 Pre-Training Data & Settings
To examine the effect of pre-training data coverage on downstream performance, we pre-trained ArTST from scratch [^2] We used the scripts and configurations provided in github.com/mbzuai-nlp/ArTST on both MSA and dialectal data. We sourced our data from various dataset: MGB2 Ali et al. (2016), QASR Mubarak et al. (2021) MGB3 Ali et al. (2017), MGB5 Ali et al. (2019), ClArTTS Kulkarni et al. (2023), ASC Halabi et al. (2016), and Common Voice Ardila et al. (2019). We also used MADAR Bouamor et al. (2018) and NADI Abdul-Mageed et al. (2023) text datasets for pre-training. In our experiments, we compare two variants of ArTST:
  • v1: This variant is as described in Toyin et al. (2023), which was pre-trained only on MSA.
  • v2: In this variant, we use a mixture of MSA and dialectal data in pre-training.
See Table 1 for details of the datasets used in pre-training the two variants.
@@
Table 2: Summary of Dataset Statistics for Fine-Tuning: Hours of Audio, Word Counts, and Associated Dialects. Multiple is mix of several dialects not neccessary from the listed dialects (no information from the source)
3.3 Fine-Tuning Data & Settings
The dataset details used for fine-tuning are shown in table 2. Note that, for MGB5 and MGB3, as the data is based on YouTube videos, many of the originally referenced videos are no longer available, so at the time of our experiments, only 2.5 hours of training were available for MGB3 and 2 hours for MGB5. Furthermore, multi dialectal datasets, such as MASC Al-Fetyani et al. (2021), have unbalanced representation of dialects. The high-resource dialects in our collection include SAU, SYR, EGY, and MSA; each has at least 200 hours of transcribed ASR data. UAE, MOR, JOR, LEB, IRA, and TUN have a medium amount of fine-tuning data between 10 and 50 hours. KUW and PAL are low-resource dialects with less than 10 hours of transcribed data in total. Finally, we left ALG, YEM, and SUD from the MASC dataset for zero-shot testing.
Figure 2 illustrates the distribution of dialectal data in pre-training and fine-tuning stages for a direct comparison. We exclude MSA from the figure as it has disproportionally more data than all dialects.
Figure 2: Distribution of dialectal speech data in pre-training and fine-tuning. MSA is not shown.
@@ 3.4 Experimental Settings
For partitioning the data into training, development, and test sets, we adhered to the official splits provided with each dataset. We followed the data preparation and training methodology established in the original ArTST implementation. For comprehensive details regarding the model architecture and data preprocessing, readers are directed to Toyin et al. (2023). The pre-training phase was executed on a cluster of 4 A100 GPUs over a duration of 21 days. We used Adam optimizer with a learning rate of $2\times 10^{-4}$, spanning 335K updates, and a warm-up phase of 64K updates. The maximum speech token length was set at 250K (equivalent to 15.625 seconds). The fine-training phase was executed on a cluster with one A100 GPUs over a duration of 7 days (MGB2, QASR, MASC) and 2 days (MGB3, MGB5, etc.). We used Adam optimizer with a tri stage scheduler for learning rate of $6\times 10^{-5}$.
@@ 4 Effect of Pre-Training Data
In this section, we examine the effect of pre-training on downstream ASR performance. As described in section 3.2, we use two variants of the model, one pre-trained mainly on MSA (v1), and one trained on additional dialectal data (v2). Note that pre-training does not utilize aligned speech and text; it incorporates un-aligned speech and text data for self-supervised learning.
@@ 4.1 Benchmarking MSA Performance
We first report results on benchmark datasets to compare the performance of both models against the state of the art. MGB2 is the main benchmark for MSA speech recognition. We evaluate the performance of both v1 and v2 models fine-tuned on MGB2 on Table 3 compared to existing SOTA models. The results confirm that incorporating dialectal data in pre-training does not negatively affect the performance on MSA.
@@
ArTST v2 (ours) 12.49 6.44
ArTST v2 (ours) + LM 12.39 6.51
Table 3: Comparing ArTST performance against models reported in Hussein et al. (2022) and Toyin et al. (2023), which include best performing model previously reports on MGB2.
@@ 4.2 Benchmarking Dialectal Performance
Tables 5 and 5 show the performance of the models on the dialectal MGB3 (Egyptian) and MGB5 (Morocooan) benchmarks. Each of these benchmarks contain multiple references as dialectal speech has no standard spelling. We report the average and multi-refrence WER for our model variants, and compare against the best model in each challenge, as well as the state-of-the-art model in each benchmark. Each model is first fine-tuned on MSA, then fine-tuned again on the target MGB train set. We also report the results of the large multilingual models: Whisper and MMS, fine-tuned on the same set. We refer to the data the model is previously adapted on as ‘Adaptation’ data. Starting with MSA data before fine-tuning on the target dialect has previously been established as an effective strategy for dialectal ASR Ali et al. (2017).
@@
Mixat 56.01 61.51 39.1 35.05 34.29 33.40
TARIC-SLU 138.14 93.54 107.56 106.46 14.70 14.80
ParallelCorp 99.17 83.16 128.72 141.92 9.57 9.31
SADA 82.16 78.28 39.41 29.77 39.24 29.91
MASC Al-Fetyani et al. (2021)
@@
Table 6: WER (%) in zero-shot and fine-tuning settings. v1 and v2 models are first fine-tuned on MGB2 (MSA).
In MGB3, dialectal pre-training (v2) results in about 4% absolute reduction in WER. Smaller improvement in terms of Average WER is observed for MGB5, where there is no clear advantage observed using v2. We speculate that this difference is a result of the pre-training having a lot more Egyptian than Moroccan data (see Figure 2).

### v3

@@ Abstract
Developing robust automatic speech recognition (ASR) systems for Arabic requires effective strategies to manage its diversity. Existing ASR systems mainly cover the modern standard Arabic (MSA) variety and few high-resource dialects, but fall short in coverage and generalization across the multitude of spoken variants. Code-switching with English and French is also common in different regions of the Arab world, which challenges the performance of monolingual Arabic models. In this work, we introduce a suite of ASR models optimized to effectively recognize multiple variants of spoken Arabic, including MSA, various dialects, and code-switching. We provide open-source pre-trained models that cover data from 17 Arabic-speaking countries, and fine-tuned MSA and dialectal ASR models that include at least 11 variants, as well as multi-lingual ASR models covering embedded languages in code-switched utterances. We evaluate ASR performance across these spoken varieties and demonstrate both coverage and performance gains compared to prior models.
@@ 1 Introduction
The advent of large self-supervised acoustic models has transformed speech technology, enabling transfer learning and improving performance for both high-resource and low-resource languages. Prominent examples of such models include various versions of wav2vec Schneider et al. (2019); Baevski et al. (2020), HuBERT Hsu et al. (2021), and SpeechT5 Ao et al. (2022), which have predominantly been trained on English datasets. Their multi-lingual variants, e.g. XLS-R Babu et al. (2022) with 53 and 128 languages, in addition to models that include both self-supervised and supervised pre-training, such as Whisper Radford et al. (2023) with approximately hundred supported languages, MMS Pratap et al. (2024) with thousands of languages, and UniSpeech Wang et al. (2021), illustrate the potential for cross-lingual transfer learning for more inclusive ASR. Yet, while these models indeed show great potential for transfer learning to new languages, even those unseen in training, they remain suboptimal for some target languages. A case in point is the Arabic Text and Speech Transformer (ArTST), a model pre-trained exclusively on Arabic, which has demonstrated superior performance for Modern Standard Arabic (MSA), surpassing larger multilingual models like Whisper and MMS in benchmark tests, in addition to establishing a new state-of-the-art performance compared to previous efforts for Arabic ASR. This highlights the advantage of monolingual pre-training when large amounts of unlabeled data for the target language are available. While the model showed some potential for dialectal coverage, it was trained and validated exclusively on MSA data, which questions its applicability for spoken dialectal variants of Arabic. Evaluations on code-switched data showed poor performance of ArTST compared to multilingual models Kadaoui et al. (2024), demonstrating the delicate trade-off between monolingual and multilingual optimization. Arabic is a pluricentric language Schuppler et al. (2024), diverse in regional variations, and models trained on MSA frequently struggle to adapt to these variations. This limitation is particularly acute given that many Arabic dialects are underrepresented and considered low-resource in speech technology research. Consequently, there is a need for optimized ASR systems that embrace, rather than overlook, the linguistic diversity of the Arabic-speaking world.
@@
In light of these challenges, we conduct various investigations aimed at understanding and enhancing the dialectal diversity and performance of Arabic ASR systems. We focus on four inquiries aimed at optimizing potential strategies for integrating dialectal variation into ASR systems. First, we measure the impact of incorporating a broad collection of Arabic dialects during the model’s pre-training phase. We hypothesize that a wider dialectal foundation could improve the model’s performance across various dialects in the fine-tuning stage. Second, we quantify the comparative effectiveness of dialect-specific fine-tuning versus a more holistic, multi-dialectal fine-tuning strategy. The third question examines the model’s capacity for zero-shot transfer to dialects not explicitly included in fine-tuning. Finally, we evaluate the model on code-switched utterances, and examine the effect of multilingual pre-training and fine-tuning on both monolingual and code-switched datasets. Our key findings from experiments spanning over 17 variants of spoken Arabic are: (1) pre-training with more data and wider dialectal coverage improves performance across most dialectal variants, including MSA, (2) multi-dialectal fine-tuning improves performance for low-resource dialects, but may not be optimal for high-resource dialects, (3) multi-dialectal pre-training and fine-tuning has higher potential for zero-shot transfer to unseen dialects, and (4) multi-lingual pre-training and fine-tuning greatly boosts performance on code-switching, while negatively impacting monolingual performance due to language interference. Our pretraining checkpoints and joint models were trained exclusively on open-source data and are released as open-source, open-weights models. All scripts required to reproduce our results including model training, evaluation, and checkpoints, are publicly available  [^1] https://github.com/mbzuai-nlp/ArTST.
@@ 2 Related Work
Arabic speech recognition research has a long history, but the majority of this research has focused on Modern Standard Arabic (MSA), the formal variant predominant in news broadcasts and official communications. A review article covering peer-reviewed publications between 2011 and 2021 estimates that 89% of papers on Arabic ASR cover MSA, and only a quarter cover some dialectal variant of Arabic Dhouib et al. (2022). Recent research in end-to-end ASR for Arabic, as presented in Hussein et al. (2022), demonstrates the potential of contemporary deep learning techniques for decoding spoken Arabic, but it was also limited to MSA. Large-scale multilingual ASR models, such as Whisper Radford et al. (2023) and MMS Pratap et al. (2024), cover many languages within their scope, including Arabic. They utilize language embeddings or adapters to enhance language coverage and performance within the same model, but their performance across languages vary considerably. For instance, while Whisper demonstrates strong zero-shot capabilities on MSA, its zero-shot accuracy drops substantially on dialects, and additional finetuning on dialectal data is needed to improve performance Waheed et al. (2023). Recent work has shown that smaller, Arabic-specific student models distilled from large models like Whisper can achieve comparable or even superior performance, especially on dialectal data, with good generalization to unseen dialects Waheed et al. (2024). Specialized Arabic models like ArTST Toyin et al. (2023), primarily pre-trained on MSA, have shown competitive results on MSA tasks and even outperformed larger multilingual models on some mixed-dialect datasets like QASR Mubarak et al. (2021). However, due to its monolingual pretraining, the model was shown to perform poorly in code-switched Arabic-English speech Al Ali and Aldarmaki (2024). This illustrates that strong MSA performance is not a reliable predictor for dialectal or code-switching capabilities, with a substantial gap persisting between SOTA MSA and dialectal performance. The development of diverse datasets such as QASR (multi-dialectal broadcast news) Mubarak et al. (2021), SADA (Saudi Arabic) Alharbi et al. (2024), ArZen (Egyptian-English code-switching) Al-Sabbagh (2024), and Mixat (Emirati-English code-switching) Al Ali and Aldarmaki (2024), and other public datasets covering various dialects and code-switched instances presents an opportunity for improving the generalization of ASR systems to diverse spoken varieties.
3 Methodology
Based on prior work, we start with the premise that monolingual training is more suitable for maximizing performance in Arabic ASR. However, the current Arabic SOTA models have limited coverage of spoken varieties and struggle with code-switching due to their monolingual training. Our objective is to maximize performance while also widening the coverage to include MSA, regional dialects, and instances of code-switching. To that end, we start with an Arabic-centric ASR model, ArTST Toyin et al. (2023), as the foundation for our investigation. Figure 1 illustrates the high-level architecture of ArTST for self-supervised pre-training and fine-tuning. This model is based on the SpeechT5 approach Ao et al. (2022), and supports multi-modal fine-tuning. The first version of the model was pre-trained on the MGB2 Ali et al. (2016) dataset, which consists of newswire data, mainly in MSA, with a small subset of dialectal variants. In this work, we attempt to understand the factors that enable both high performance and wide coverage; we explore the following questions:
@@ 3 Methodology
    Is pretraining on dialectal data beneficial for improving down-stream dialectal performance, and would it negatively impact MSA performance?
@@
    Is it better to finetune ASR models jointly on multiple dialects or fine-tune on a specific target dialect?
@@
    Can we optimize performance in code-switched utterances using multilingual pretraining?
  5. 5.
    What is the effect of multilingual pretraining and finetuning on monolingual Arabic performance? (i.e. language interference).
The remaining sections detail our experimental settings and findings of these questions.
@@ 3.2 Pre-Training Data & Settings
To examine the effect of pre-training data coverage on downstream performance, we pre-trained ArTST from scratch [^2] We used the scripts and configurations provided in github.com/mbzuai-nlp/ArTST on both MSA and dialectal data. We sourced our data from various datasets, including: MGB2 Ali et al. (2016), QASR Mubarak et al. (2021) MGB3 Ali et al. (2017), MGB5 Ali et al. (2019), ClArTTS Kulkarni et al. (2023), ASC Halabi et al. (2016), and Common Voice Ardila et al. (2020), SADA Alharbi et al. (2024), and others. We also used MADAR Bouamor et al. (2018) and NADI Abdul-Mageed et al. (2023) text datasets for pretraining. In our experiments, we compare the following:
  • v1: This variant is as described in Toyin et al. (2023), pretrained only on MSA.
  • v2: In this variant, we use a mixture of MSA and dialectal data in pretraining.
  • v3: In this variant, we use a mixture of MSA, dialectal, and multilingual data in pretraining.
See Table 12 in Appendix A for details of all the datasets used in pre-training.
@@
Table 1: Summary of Dataset Statistics for Fine-Tuning: Hours of Audio, Word Counts, and Associated Dialects. Multiple is mix of several dialects not neccessary from the listed dialects (no information from the source).
3.3 Dialectal Fine-Tuning
The datasets we use for dialectal fine-tuning are shown in Table 1. We collected as many open-source data as needed to maximize coverage of dialects. Note that, for MGB5 and MGB3, as the data is based on YouTube videos, many of the originally referenced videos are no longer available, so at the time of our experiments, only 2.5 hours of training were available for MGB3 and 2 hours for MGB5. Furthermore, multi dialectal datasets, such as MASC Al-Fetyani et al. (2021), have unbalanced representation of dialects. The high-resource dialects in our collection include SAU, SYR, EGY, and MSA; each has at least 200 hours of transcribed ASR data. UAE, MOR, JOR, LEB, IRA, and TUN have a medium amount of fine-tuning data between 10 and 50 hours. KUW and PAL are low-resource dialects with less than 10 hours of transcribed data in total. Finally, we left ALG, YEM, and SUD from the MASC dataset for zero-shot testing.
Figure 2 illustrates the distribution of dialectal data we use for pre-training and fine-tuning our dialectal model. We exclude MSA from the figure as it has disproportionally more data than all dialects.
3.4 Multi-Lingual Fine-Tuning
In addition to the above dialectal data, we used English, French, and Spanish sets for multi-lingual fine-tuning and code-switching experiments described in section 8. English and French are commonly spoken in various Arabic-speaking countries, and to a lesser extent, Spanish is spoken in some parts of North Africa. More details about the datasets used in these experiments are provided in section 8.
@@ 3.5 Experimental Settings
For partitioning the data into training, development, and test sets, we adhered to the official splits provided with each dataset. We followed the data preparation and training methodology established in the original ArTST implementation. For comprehensive details regarding the model architecture and data preprocessing, readers are directed to Toyin et al. (2023).
Figure 2: Distribution of dialectal speech data in pre-training and fine-tuning. MSA data are not shown.
Computational Details
The pre-training was executed on a cluster of 4 A100 GPUs over a duration of 14 to 21 days for each model. We used Adam optimizer with a learning rate of $2\times 10^{-4}$, spanning 335K updates, and a warm-up phase of 64K updates. The maximum speech token length was set at 250K (equivalent to 15.625 seconds). Each fine-tuning experiment was run on one A100 GPU over a duration of 7 days (MGB2, QASR, MASC, SADA) or 2 days for smaller sets (MGB3, MGB5, etc.). We used Adam optimizer with a tri-stage scheduler with learning rate of $6\times 10^{-5}$. The total computational budget for all experiments is estimated to be $\sim$6000 GPU-hours.
@@ 4 Effect of Pre-Training Data
We first examine the effect of pre-training on downstream ASR performance. As described in section 3.2, we compare a model pre-trained mainly on MSA (ArTST v1), and our multi-dialectal version (henceforth v2). Note that pre-training does not utilize aligned speech and text; it incorporates unaligned speech and text data for self-supervised learning. For these experiments, we use the same finetuning data, and only vary the pretraining sets.
@@ 4.1 Benchmarking MSA Performance
We first report results on benchmark datasets to compare the performance of both models against the state-of-the-art. MGB2 is the main benchmark for MSA speech recognition. We evaluated the performance of ArTST v1 and v2 fine-tuned in MGB2, compared to existing SOTA models, in Table 2. The results show that incorporating dialectal data in pretraining does not negatively affect MSA performance, as we achieve the best WER of 12.48% and 12.39%, with and without LM fusion.
@@
v2 12.49 6.44
v2 + LM 12.39 6.51
Table 2: Comparing our performance against models reported in Hussein et al. (2022) and Toyin et al. (2023), which include the best performing models previously reported on MGB2.
@@ 4.2 Benchmarking Dialectal Performance
Tables 4 and 4 show the performance of the models on the dialectal MGB3 (Egyptian) and MGB5 (Moroccan) benchmarks. Each of these benchmarks contain multiple references as dialectal speech has no standard spelling. We report the average and multi-refrence WER for our model variants, and compare against the best model in each challenge, as well as the SOTA model in each benchmark. Each model is first fine-tuned on MSA, then fine-tuned again on the target MGB train sets. We also report the results of the large multilingual models: Whisper Radford et al. (2023) and MMS Pratap et al. (2024), fine-tuned on the same set. We refer to the Arabic data the models are previously fine-tuned on as ‘Adaptation’ data. Starting with MSA data before fine-tuning on the target dialect has previously been established as an effective strategy for dialectal ASR Ali et al. (2017).
@@
TARIC-SLU    (TUN) 138.14 93.54 107.56 106.46 14.70 14.80
ParallelCorp (MULT) 99.17 83.16 128.72 141.92 9.57 9.31
SADA    (SAU) 82.16 78.28 39.41 29.77 39.24 29.91
MASC
@@
Table 5: WER (%) in zero-shot and fine-tuning settings. We compare zero-shot performance of Whisper, MMS, ArTST v1, and Our dialectal pretraining (v2). ArTST v1 and v2 are finetuned on MGB2 (MSA), whereas Whisper and MMS are finetuned/pretrained with multi-lingual data, including Arabic.
In MGB3, dialectal pretraining (v2) results in about 4% absolute reduction in WER, establishing a new SOTA result on this benchmark. Smaller improvement in terms of Average WER is observed for MGB5, where there is no clear advantage observed using our dialectal version. This difference is likely a result of our pre-training having a lot more Egyptian than Moroccan data (see Figure 2).